OpenStudy (anonymous):

Use the definition of the meter to determine how far light travels in 1ns.

OpenStudy (anonymous):

By definition: http://www.bipm.org/en/CGPM/db/17/1/ light travels one metre in 1/299 792 458 seconds. Now s=v*t and therefore v=(1m/1/299 792 458 s)=299 792 458 m/s. Now, multiplying it my 10^-9 s (one nanosecond) we get 0.299792458 metres.

OpenStudy (anonymous):

this is a conversion problem.. start with the speed of light in meters...3 X 10 ^8 m/s x 1s/10^9 ns.. do the math then. In conversion make sure you cancel out units. You should be left with m/ns 3 X 10^-1 m/ns

OpenStudy (anonymous):

thank you both so much!!